Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability question

Expert replies
by Matt@VeritasPrep » Sat Aug 03, 2013 2:08 pm
The new deck has 60 cards, of which (assuming both decks are standard), 4 ≤ (# of kings) ≤ 6. (There are the four normal kings, to which she could have added one black king or two black kings.)

There are two red aces in the new deck, so the odds of picking a red ace are 2/60, or 1/30.

Let's call x the number of kings. We have

(Odds of Red Ace) + (Odds of King) > 1/8

(We can add these without adjustments, as they're mutually exclusive: when you pick a card it can't be both a red ace and a king.)

1/30 + x/60 > 1/8

(x+2)/60 > 1/8

x > 5.5, meaning we must have 6 kings.

So the deck has 60 cards, 6 of which are kings, 4 of which are black.

(Odds of Black King) + (Odds of Red Jack) = 4/60 + 2/60 = 1/15 + 1/30 = 1/10

The OA (as posted) is incorrect.
Join the discussion
Source: — Problem Solving |